# Peter van Hintum

### v1

@@ 1 Introduction
Establishing discrete counterparts to fundamental results from convex geometry and analysis is an active area of research, e.g. John’s theorem [26, 40, 41, 2, 20], Klartag and Lehec’s Slicing Theorem (formerly Bourgain’s Slicing Conjecture)[29, 1, 36, 11, 12], and the Brunn-Minkowski inequality [4, 15, 34, 9, 5, 33, 16, 21] and its stability [13, 22, 23]. The reverse direction seems to have received less attention [19]. In this note, we prove discrete versions of the Prékopa-Leindler and Borell-Brascamp-Lieb inequalities (Theorem 1.2) which can be seen as the functional versions of the recently established [21] discrete Brunn-Minkowski inequality as conjectured by Ruzsa [38].
@@ 1.2 Discrete Inequalities
Some discrete version of the Prékopa-Leindler inequality have been proved previously in the style of the Ahlswede–Daykin inequality (or Four Function Theorem) [28, 17, 24]. Strong as these results are, the inequality presented here stays closer to the spirit of the Prékopa-Leindler inequality. In [25], Iglesias and Yepes establish a Borell-Brascamp-Lieb inequality in $\mathbb{Z}^{n}$ providing a lower bound based on removing the largest hyperplane sections of $f$.
@@ 1.3 From Brunn-Minkowski to Borell-Brascamp-Lieb
There are many proofs of the Borell-Brascamp-Lieb inequality and even more of the Prékopa-Leindler inequality. The approach presented here provides a general method to use Brunn-Minkowski inequalities to prove Borell-Brascamp-Lieb inequalities. Though not explicitly stated as such this approach has been present in this field of study for some time.
@@
A simple computation (see e.g. [14, Lemma 4.5]) shows that
@@ 1.4.1 Gaussian measure
Recall that a function $f\colon\mathbb{R}^{d}\to\mathbb{R}_{\geq 0}$ is even if f(x)=f(x)f(x)=f(-x) and quasiconcave if all its level sets $F_{t}=\{x\in\mathbb{R}^{d}:f(x)>t\}$ are convex. In particular any log-concave function is quasiconcave. Using Proposition 1.4, Theorem 1.8 has the following corollary.
@@ 1.4.4 Kemperman’s theorem
Kemperman’s theorem [27] gives a very general, albeit from the perspective of the Brunn-Minkowski inequality weak, lower bound on the size of sumsets in a very general class of groups. Let $G$ be a unimodular, locally compact group with Haar measure $\mu$, then it says the following.
@@ 1.4.6 Riemannian Manifolds
This example differs from the others as the result that was originally proved was a Borell-Brascamp-Lieb type inequality [7], but it is included anyway for illustration. There has been extensive study of geometric inequalities on manifolds. For the sake of brevity, I will refer to the original paper by Cordero-Erausquin, McCann, and Schmuckenschläger [7] for precise definitions and conditions. Given a sufficiently well-behaved manifold $\mathcal{M}$ with geodesic distance $d$, define the following averaging operation. For $x,y\in\mathcal{M}$ let $S_{\lambda}(x,y):=\{m\in\mathcal{M}:d(x,m)=\lambda d(x,y)\text{ and }d(m,y)=(1-\lambda)d(x,y)\}$, the collection of midpoints between $x$ and $y$. For subsets $A,B\subset\mathcal{M}$, let $S_{\lambda}(A,B):=\bigcup_{x\in X,y\in Y}S_{\lambda}(x,y)$. The size of $S_{\lambda}(A,B)$ one might expect depends on the curvature of $\mathcal{M}$ and the distance between the sets. The paper gives precise descriptions of this effect, but let me just give the following simple consequence.
@@
Let $\mathcal{M}$ an $d$-dimensional manifold with non-negative curvature with measure $\mu$ and $\lambda\in(0,1)$. Let $A,B\subset\mathcal{M}$ measurable so that $S_{\lambda}(A,B)$ is also measurable, then

### v2

@@ 1 Introduction
Establishing discrete counterparts to fundamental results from convex geometry and analysis is an active area of research, e.g. John’s theorem [30, 47, 48, 3, 22], Klartag and Lehec’s Slicing Theorem (formerly Bourgain’s Slicing Conjecture)[33, 1, 42, 14, 15], and the Brunn-Minkowski inequality [5, 18, 40, 12, 6, 29, 39, 19, 24] and its stability [16, 25, 26]. The reverse direction seems to have received less attention [23]. In this note, we prove discrete versions of the Prékopa-Leindler and Borell-Brascamp-Lieb inequalities (Theorem 1.2) which can be seen as the functional versions of the recently established [24] discrete Brunn-Minkowski inequality as conjectured by Ruzsa [44].
@@ 1.2 Discrete Inequalities
Some discrete version of the Prékopa-Leindler inequality have been proved previously in the style of the Ahlswede–Daykin inequality (or Four Function Theorem) [32, 20, 27]. Others use rounding to get strong enough bounds on $h$ to prove Prékopa-Leindler type inequalities [29, 37]. Strong as these results are, the inequality presented here stays closer to the spirit of the Prékopa-Leindler inequality. In [28], Iglesias and Yepes establish a Borell-Brascamp-Lieb inequality in $\mathbb{Z}^{n}$ providing a lower bound based on removing the largest hyperplane sections of $f$.
@@ 1.3 From Brunn-Minkowski to Borell-Brascamp-Lieb
There are many proofs of the Borell-Brascamp-Lieb inequality and even more of the Prékopa-Leindler inequality. The approach presented here provides a general method to use Brunn-Minkowski inequalities to prove Borell-Brascamp-Lieb inequalities. The basic idea is to apply the Brunn-Minkowski inequality to pairs of level sets of the functions in question and to use an optimal transport map to find an appropriate pairing. This approach has been present in the field for some time in varying degrees of explicitness. Recently, Malliaris, Melbourne, Roberto, and Roysdon [38] explored the connection between these inequalities through an approach similar to Proposition 1.4 below. In a short note, Cordero-Erausquin [8] clarified this connection further and provided some historical notes and simple proofs.
@@
A simple computation (see e.g. [17, Lemma 4.5]) shows that
@@ 1.4.1 Gaussian measure
Recall that a function $f\colon\mathbb{R}^{d}\to\mathbb{R}_{\geq 0}$ is even if f(x)=f(x)f(x)=f(-x) and quasiconcave if all its level sets $F_{t}=\{x\in\mathbb{R}^{d}:f(x)>t\}$ are convex. In particular any log-concave function is quasiconcave. Using Proposition 1.4, Theorem 1.8 has the following corollary (also obtained in e.g. [38, 9, 2]).
@@ 1.4.3 Unconditional measures
This corollary in fact also provides the induction step for an inductive proof of Theorem 1.14 as follows.
Proof of Theorem 1.14 and 1.15.
Proceed by induction on dimension $d$. For $d=1$, Theorem 1.14 is trivial and 1.15 follows from Proposition 1.4 and 1.5. For dimension $d$, consider the function $f:\mathbb{R}^{d-1}\to\mathbb{R}_{\geq 0},x\mapsto\mathcal{H}^{1}(A\cap(\{x\}\times\mathbb{R}))$ which measures the length of the fibre of $A$ above $x$, so that $\int_{\mathbb{R}^{d-1}}f(x)dx=|A|$. Analogously define $g,h$ for the sets $B$ and $A^{\lambda}B^{1-\lambda}$. Note that as all sets are unconditional, these functions satisfy $h(x^{\lambda}y^{1-\lambda})\geq f(x)^{\lambda}g(y)^{1-\lambda}$ for all $x,y\in\mathbb{R}^{d-1}$. Hence, by 1.15 for dimension d1d-1, we find $|A^{\lambda}B^{1-\lambda}|=\int_{\mathbb{R}^{d-1}}h(x)dx\geq\int_{\mathbb{R}^{d-1}}f(x)dx=|A|$, so that Theorem 1.14 holds for dimension $d$. 1.15 for dimension $d$ follows by Proposition 1.4 and 1.5 concluding the induction. ∎
@@ 1.4.4 Kemperman’s theorem
Kemperman’s theorem [31] gives a very general, albeit from the perspective of the Brunn-Minkowski inequality weak, lower bound on the size of sumsets in a very general class of groups. Let $G$ be a unimodular, locally compact, connected111[46] considers a relaxation of the connectedness condition, which in turn has its own functional corollary. group with Haar measure $\mu$, then it says the following.
@@ 1.4.6 Riemannian Manifolds
This example differs from the others as the result that was originally proved was a Borell-Brascamp-Lieb type inequality [10], but it is included anyway for illustration. There has been extensive study of geometric inequalities on manifolds. For the sake of brevity, I will refer to the original paper by Cordero-Erausquin, McCann, and Schmuckenschläger [10] for precise definitions and conditions. Given a sufficiently well-behaved manifold $\mathcal{M}$ with geodesic distance $d$, define the following averaging operation. For $x,y\in\mathcal{M}$ let $S_{\lambda}(x,y):=\{m\in\mathcal{M}:d(x,m)=\lambda d(x,y)\text{ and }d(m,y)=(1-\lambda)d(x,y)\}$, the collection of midpoints between $x$ and $y$. For subsets $A,B\subset\mathcal{M}$, let $S_{\lambda}(A,B):=\bigcup_{x\in X,y\in Y}S_{\lambda}(x,y)$. One might expect the size of $S_{\lambda}(A,B)$ to depend on the curvature of $\mathcal{M}$ and the distance between the sets. The paper gives precise descriptions of this effect, but let me just give the following simple consequence.
@@
Let $\mathcal{M}$ an $d$-dimensional manifold with non-negative curvature with measure $\mu$, and $\lambda\in(0,1)$. Let $A,B\subset\mathcal{M}$ measurable so that $S_{\lambda}(A,B)$ is also measurable, then
@@ 2 Proof of the discrete Borell-Brascamp-Lieb inequality
Acknowledgments
The author is grateful to Gautam Aishwarya and Francisco Marin Sola for pointing me to [38] and [8]. The author would like to additionally thank Takashi Satomi for pointing out the missing connectedness condition in Kemperman’s theorem and its corollary.